OpenStudy (anonymous):

how do I figure out that csc^2x*sinx-cotx*sinx*tanx=cosx*cotx

OpenStudy (jdoe0001):

\(\bf csc^2(x)sin(x)-cot(x)sin(x)tan(x)=cos(x)cot(x) \\ \quad \\ \implies \cfrac{1}{sin^2(x)}\cdot sin(x)-\cfrac{cos(x)}{sin(x)}\cdot sin(x)\cdot \cfrac{sin(x)}{cos(x)}\) what would that left-side give you?